OpenStudy (anonymous):

If A(2, 3) and B(7, 15), find AB. (Remember: AB means "the distance between points A and B.")

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 2}}\quad ,&{\color{blue}{ 3}})\quad B&({\color{red}{ 7}}\quad ,&{\color{blue}{ 15}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (jdoe0001):

hmm scratch that lemme get the "other" formula =)

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 2}}\quad ,&{\color{blue}{ 3}})\quad B&({\color{red}{ 7}}\quad ,&{\color{blue}{ 15}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)